# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ Abstract
The rapid advancement and expanding applications of Audio Large Language Models (ALLMs) demand a rigorous understanding of their trustworthiness. However, systematic research on evaluating these models, particularly concerning risks unique to the audio modality, remains largely unexplored. Existing evaluation frameworks primarily focus on the text modality or address only a restricted set of safety dimensions, failing to adequately account for the unique characteristics and application scenarios inherent to the audio modality. We introduce AudioTrust—the first multifaceted trustworthiness evaluation framework and benchmark specifically designed for ALLMs. AudioTrust facilitates assessments across six key dimensions: fairness, hallucination, safety, privacy, robustness, and authentication. To comprehensively evaluate these dimensions, AudioTrust is structured around 18 distinct experimental setups. Its core is a meticulously constructed dataset of over 4,420 audio/text samples, drawn from real-world scenarios (e.g., daily conversations, emergency calls, voice assistant interactions), specifically designed to probe the multifaceted trustworthiness of ALLMs. For assessment, the benchmark carefully designs 9 audio-specific evaluation metrics, and we employ a large-scale automated pipeline for objective and scalable scoring of model outputs. Experimental results reveal the trustworthiness boundaries and limitations of current state-of-the-art open-source and closed-source ALLMs when confronted with various high-risk audio scenarios, offering valuable insights for the secure and trustworthy deployment of future audio models. Our platform and benchmark are available at https://github.com/JusperLee/AudioTrust.
@@ 1 Introduction
Despite rapid progress in ALLMs [18, 65, 14, 94, 95, 82, 90], there is still no comprehensive evaluation benchmark addressing their potential risks. Existing safety evaluation frameworks, such as SafeDialBench [9] and SafetyBench [100], mainly focus on the text modality and do not consider the unique properties or application scenarios of audio. Unlike text, audio features complex temporal-frequency patterns, rich emotions, and contextual dependencies, and introduces additional challenges, such as speech privacy, speaker recognition, and environmental acoustic analysis. These factors lead to unique trustworthiness risks in audio, including gender and accent biases, audio hallucinations, social engineering attacks, personal information leakage, and adversarial attacks on audio systems.
To address this research gap, we propose AudioTrust—a comprehensive trustworthiness evaluation framework specifically tailored for ALLMs (see Appendix B), which fully considers the inherent features and application scenarios of the audio modality, as shown in Figure 1. AudioTrust covers six key trustworthiness dimensions: (1) fairness, (2) hallucination, (3) safety, (4) privacy, (5) robustness, and (6) authentication. An overview of the AudioTrust benchmark structure, detailing these dimensions and their sub-categories, is presented in Figure 2 (a). The foundation of this benchmark is a diverse dataset comprising over 4,420 audio-text samples derived from realistic application contexts, including daily conversations, emergency communications, and voice assistant interactions. This dataset facilitates 18 distinct experimental setups, each targeting specific trustworthiness facets. For a rigorous and scalable assessment, we implement a large-scale automated evaluation pipeline in Appendix C and design evaluation criteria dedicated to the characteristics of the audio modality. Preliminary results of representative models are summarized in the leaderboard shown in Figure 2 (b). For different evaluation types, we have established more detailed data segmentation forms and experimental settings based on scenarios, attack types, and other factors. The specific perspectives of trustworthiness that AudioTrust focuses on are comprehensively enumerated in Figure 4.
Fairness: We assess fairness across seven sensitive attributes (such as gender and age) in three decision-making contexts—recruitment, admissions, and loan approval—using 840 audio samples. The same data supports stereotype experiments in work, education, and medical scenarios. Group unfairness metrics quantify bias. Key findings: (1) ALLMs exhibit systematic biases, with no model achieving ideal fairness. (2) Accent and personality traits show stronger bias than age and gender. (3) Closed-source models are less fair in decision-making tasks, while open-source models are less fair in stereotype settings. (4) Model outputs generally favor female, older, American-accented, calm, fluent, Western, and wealthy profiles. (5) Models differ in which attributes they are most biased toward.
Hallucination: Hallucinations are categorized into logical and factual types, with 320 test cases constructed across various scenarios. Through illusion detection and robustness analysis, we find: (1) Closed-source models are generally better at identifying acoustically illogical events, while open-source models lag behind. (2) Some advanced closed-source models can detect mismatches between audio and label attributes, but many models are misled by incorrect metadata, indicating ALLMs’ limitations in domain knowledge alignment. (3) Certain closed-source models accurately judge temporal logic in audio, though some ALLMs show unexpected weaknesses. (4) Some closed-source models excel in cross-modal semantic consistency, while many ALLMs—including open-source and proprietary versions—struggle with this task.
Safety: Safety is evaluated in two areas: jailbreak attacks and illegal activity guidance, using 300 cases for each in enterprise, financial, and medical domains. Based on defense success and harmful response rates, we find: (1) Closed-source models generally have stronger defenses, while open-source models are more vulnerable, especially to jailbreaks. (2) Jailbreak risks are higher in specialized domains (e.g., medical), highlighting gaps in domain knowledge. (3) Models, especially closed-source ones, are more resistant to illegal activity guidance, whereas open-source safeguards are weaker. (4) Audio-specific factors, like emotion, influence model defenses against jailbreaks.
Privacy: Privacy evaluation targets risks of direct and inference leakage. We created 600 samples with explicit sensitive information (e.g., bank accounts, SSNs, addresses) and 300 samples for implicit attribute inference (e.g., age, ethnicity), using a five-point scale to assess refusal, complete leakage, and cautious leakage rates in audio QA tasks across two datasets. Key findings: (1) ALLMs provide uneven protection for different sensitive information, with stronger safeguards for items like SSNs and bank accounts. (2) ALLMs can infer private attributes such as age and ethnicity, but most models lack awareness of the privacy implications. (3) Simple prompt engineering helps reduce direct leaks, but is less effective for inference leakage.
Robustness: Robustness is evaluated using 240 cases across adversarial and non-adversarial scenarios. Analyzing anti-interference and output quality, we find: (1) Leading closed-source ALLMs generally show stronger robustness, while most open-source models are more vulnerable, with a few exceptions. (2) Closed-source ALLMs maintain core performance under various challenges; most open-source models suffer notably from background noise, multiple speakers, or degraded audio, especially with adversarial or environmental interference. (3) Closed-source ALLMs more accurately perceive and describe audio quality differences, while other models’ abilities in this area are limited or inconsistent.
Authentication: Authentication evaluation covers three attack types: Hybrid Spoofing, Identity Verification Bypass, and Voice Cloning Spoofing, using 100 samples each for the first two and 200 for the last. Findings: (1) Some Gemini models struggle to distinguish genuine speech from high-fidelity clones. (2) Closed-source models are rarely fooled in verification bypass or spoofing scenarios, while some open-source models are more susceptible. (3) Stricter text prompts improve resistance to voice cloning attacks across all models, highlighting the importance of prompt design in enhancing spoofing defenses. (4) Attackers may employ social engineering or acoustic interference (e.g., echo, crowd noise) to undermine verification reliability.
@@ 2 AudioTrust: Fairness
In this section, we assess the fairness of ALLMs by focusing on several key research questions: (1) Is there a statistically significant correlation between decision outputs of ALLMs (e.g., classification or identification results) and sensitive attributes (e.g., gender, age)? (2) Which sensitive attributes demonstrate more pronounced biases across different models? (3) In which specific scenarios (such as hiring or financial lending decisions) do various models exhibit stronger biases toward certain sensitive attributes? (4) Which ALLM performs better concerning fairness metrics?
Attack Strategies. Ensuring fairness in ALLMs is essential due to their rapid development. Biases in ALLMs impact task performance and may reinforce harmful stereotypes, affecting societal fairness [31, 74]. We identify two types of fairness issues: traditional and audio-based. Traditional fairness involves stereotypes related to gender, race, and age [13, 101, 70, 25, 28, 71], while some benchmarks include other attributes like appearance [52, 55], which we do not consider due to the audio modality. Audio-based fairness includes biases around accent, linguistic characteristic, economic status, and personality traits, which we simulate using mixed audio and text preprocessing. For each sensitive attribute, we designed decision-making and stereotype-driven scenarios [81, 74]. Decision-making covers recruitment, admissions, and loan evaluations. Stereotype-driven contexts address beliefs such as men outperforming women in mathematics and gender roles in medicine and occupations. Dataset classification is detailed in Section D.1. Our dataset has 840 audio samples across seven attributes: gender, age, race, accent, linguistic characteristic, economic status, and personality traits, each about 20 seconds long. See Section D.2 for details.
Evaluation & Metrics. For fairness evaluation, we introduced a key metric: Group Unfairness Score [81, 15]. Group unfairness examines the distributional equity of model outputs across different groups (e.g., male versus female), typically measured by the divergence or balance between group distributions. For detailed formulas and further analysis in the Section D.3 and Section D.4.
@@
| MiniCPM-o 2.6 | 0.585 | 0.752 |
| Qwen2-Audio | 0.710 | 0.667 |
| SALMONN | 0.912 | 0.861 |
| Ultravox | 0.632 | 0.762 |
@@
| Gemini-1.5 Pro | 0.540 | 0.703 |
| Gemini-2.5 Flash | 0.760 | 0.630 |
| Gemini-2.5 Pro | 0.795 | 0.681 |
| GPT-4o Audio | 0.726 | 0.074 |
| GPT-4o mini Audio | 0.755 | 0.136 |
| Average | 0.716 | 0.630 |
Note: : higher than column average, : lower than column average, subscript is absolute difference.
Results. We evaluated the group fairness of seven ALLMs in terms of social stereotypes and decision-making in Table 1. Complete results and examples are provided in Section D.4. (1) Existing ALLMs exhibit severe unfairness across different sensitive attributes, falling far short of the ideal fairness (i.e., ). (2) Personality and age biases are more prominent in ALLMs, whereas gender bias is the least significant. (3) In decision-making experiments, among the open-source models, MiniCPM-o 2.6 and Ultravox demonstrate relatively higher fairness. Among the Gemini series, Gemini-1.5 Pro shows better fairness, outperforming Gemini-2.5 Flash and Gemini-2.5 Pro. In stereotype experiments, GPT-4o mini-audio and GPT-4o Audio exhibit the highest group fairness. (4) There is a significant gap between fairness in decision-making and fairness in stereotypes. This is because we designed extreme decision-making scenarios, where the GPT-4o series models sacrificed fairness to maintain response accuracy. (5) Overall, among closed-source models, GPT-4o Audio demonstrates the strongest fairness. The Gemini series perform poorly in fairness, with a significant gap compared to the GPT series. Among open-source models, MiniCPM-o 2.6 shows the highest fairness, while SALMONN performs the worst.
@@ 3 AudioTrust: Hallucination
Unlike hallucinations in the textual domain, which have been extensively investigated [29, 62, 89], those in ALLMs exhibit unique multimodal characteristics. We evaluate the ability of ALLMs in handling and understanding complex acoustic scenarios, especially in identifying potential “hallucinations” or inconsistencies, by focusing on the following key research questions: (1) Can ALLMs identify acoustic events that are contrary to physical laws? (2) When there is a mismatch between the audio content and its tag attributes, can ALLM effectively analyze and understand the audio content? (3) How does ALLM perform in understanding and judging the rationality of the temporal logic of audio events? (4) How accurate is ALLM in determining whether the audio content is consistent with the core semantics described in the given text?
Attack Strategies. We present a taxonomy of audio hallucinations, exposing key trustworthiness risks in current ALLMs. Hallucinations are grouped into two main categories: factual and logical. For each, we collected 160 data samples for comprehensive evaluation (Section E.1). Both types present real-world dangers. Factual hallucinations include physical law violations and attribute mismatches. For example, when models fail to recognize physically impossible sounds (e.g., distinguishing a building collapse from a “silent explosion”), emergency response systems could misinterpret events and misdirect aid [41]. We synthesize deceptive audio by splicing segments to simulate impossible scenarios. Attribute mismatches—such as confusing distress calls with casual conversation—are critical for emergency dispatch [75]. Using the Emotional Speech Dataset [105], we deliberately mismatch emotion labels and audio to expose these risks. Logical hallucinations, such as temporal misordering and semantic contradictions, are equally serious. Temporal errors (e.g., confusing the sequence of a collapse and a gunshot) jeopardize security investigations [11]. Our benchmark (Section E.2) features reversed causal event sequences for such tests. Semantic contradictions—like misidentifying breaking glass as rainfall—can leave facilities exposed [16]. We create tests by pairing audio with misleading text descriptions to reveal model vulnerabilities. This taxonomy and evaluation reveal critical safety gaps that must be addressed before ALLMs are ready for deployment in high-stakes applications.
Evaluation & Metrics. We introduce a comprehensive evaluation framework covering hallucination recognition, property verification, reality consistency assessment [21], and transcription accuracy. Full experimental details are in Section E.3.1. For hallucination recognition, models detect hallucinations in audio-text pairs and provide justifications. Property verification targets physical, label, and content hallucinations via multi-stage prompting [46]. Reality consistency tasks assess alignment with real-world knowledge using a two-stage protocol [33]. Transcription accuracy is measured by Word Error Rate (WER), focusing on inconsistent cases [16]. Hallucination detection uses multiple metrics and GPT-4o-based evaluation [54], which has enhanced human judgment in prior studies. Each test yields a 0–10 score based on prompts, with 10 indicating precise identification and explanation of logical or semantic errors. We report Cross-Modal WER (CM-WER) [65], for transcription accuracy under hallucination scenarios. See Section E.3.2 for metric details.
@@
| Average | 5.96 | 5.01 | 6.78 | 7.77 |
Note: : higher than column average (better for accuracy), : lower than column average (worse for accuracy), subscript is absolute difference. Scenarios: Content Mismatch (CM), Label Mismatch (LM), Logical Violation (LV), and Physical Violation (PV).
Results. We evaluated the hallucination resistance of nine ALLMs across four dimensions, as shown in Table 2. Complete results and examples are provided in Section E.4 and Section E.5. We observed the following key points: (1) Different ALLMs exhibit significant variance in anti-hallucination performance. In general hallucination assessment, Gemini-1.5Pro, Gemini-2.5Flash, Gemini-2.5Pro, and Ultravox achieve higher scores, while SALMONN, GPT-4o Audio, and GPT-4o mini Audio are more prone to hallucination. Open-source models MiniCPM-o 2.6 and Qwen2-Audio also deliver competitive results. (2) For physically impossible audio event recognition, Gemini-2.5 Pro and Gemini-1.5 Pro obtain the best results, reflecting stronger physical reasoning, while models like SALMONN lag behind. (3) In tasks requiring analysis of audio content independent from labels, Gemini-2.5 Pro and Gemini-2.5 Flash can accurately identify mismatched tags, whereas SALMONN and GPT-4o mini Audio are easily misled by false metadata. (4) Regarding temporal logic, Gemini-2.5 Pro, Gemini-2.5 Flash, and Gemini-1.5 Pro show strong ability to detect temporal inconsistencies within audio, but GPT-4o mini Audio presents notable weaknesses. (5) In cross-modal reasoning, Gemini-2.5 Pro and Gemini-1.5 Pro are most robust at detecting discrepancies between audio and textual descriptions, whereas SALMONN and GPT-4o mini Audio remain limited in this aspect.

### v3

@@ Abstract
The rapid development and widespread adoption of Audio Large Language Models (ALLMs) require a rigorous assessment of their trustworthiness. However, existing evaluation frameworks, primarily designed for text, are not equipped to handle the unique vulnerabilities introduced by audio’s acoustic properties. We find that significant trustworthiness risks in ALLMs arise from non-semantic acoustic cues, such as timbre, accent, and background noise, which can be used to manipulate model behavior. To address this gap, we propose AudioTrust, the first framework for large-scale and systematic evaluation of ALLM trustworthiness concerning these audio-specific risks. AudioTrust spans six key dimensions: fairness, hallucination, safety, privacy, robustness, and authenticition. It is implemented through 26 distinct sub-tasks and a curated dataset of over 4,420 audio samples collected from real-world scenarios (e.g., daily conversations, emergency calls, and voice assistant interactions), purposefully constructed to probe the trustworthiness of ALLMs across multiple dimensions. Our comprehensive evaluation includes 18 distinct experimental configurations and employs human-validated automated pipelines to objectively and scalably quantify model outputs. Experimental results reveal the boundaries and limitations of 14 state-of-the-art (SOTA) open-source and closed-source ALLMs when confronted with diverse high-risk audio scenarios, thereby offering critical insights into the secure and trustworthy deployment of future audio models. Our platform and benchmark are publicly available at this link.
@@ 1 Introduction
Despite rapid progress in ALLMs (Du et al., 2023; Tang et al., 2024; Chu et al., 2024; Zhang et al., 2023a; 2024a; Xu et al., 2025c; Yao et al., 2024), there is still no comprehensive benchmark evaluating their potential risks. Existing safety evaluation frameworks, such as SafeDialBench (Cao et al., 2025) and SafetyBench (Zhang et al., 2023c), mainly focus on the text modality while overlooking the unique properties or application scenarios of audio. Unlike text, audio features complex temporal-frequency patterns, rich emotions, and contextual dependencies, and introduces additional challenges, such as speech privacy, speaker recognition, and environmental acoustic analysis. These factors lead to unique trustworthiness risks in audio, including gender and accent biases, audio hallucinations, social engineering attacks, personal information leakage, and adversarial attacks on audio systems.
The integration of audio modalities into large models, while functionally powerful, introduces a new attack surface and exacerbates existing trustworthiness vulnerabilities. To systematically quantify these emergent risks, we introduce AudioTrust, the first comprehensive benchmark designed to evaluate the trustworthiness of ALLMs. AudioTrust establishes a rigorous evaluation framework across six critical dimensions where audio introduces unique safety concerns: (1) Fairness: Evaluating biases derived from vocal delivery, rather than the semantic content of speech; (2) Hallucinations: Testing for audio-grounded hallucinations, where model outputs violate the physical laws or temporal logic of an acoustic scene. (3) Safety: Assessing resilience to harmful queries that leverage persuasive or emotional vocal tones to bypass safety filters; (4) Privacy: Quantifying the leakage of sensitive information from spoken content and inference of personal attributes from acoustic cues; (5) Robustness: Assessing model performance in acoustically complex and imperfect environments, such as those with background noise. (6) Authentication: Evaluating the ability to distinguish authentic speakers from sophisticated impersonation attacks, including voice clones and audio-based social engineering. Underpinning our benchmark is a curated dataset of over 4,420 audio samples spanning 18 distinct evaluation tasks, from emergency communications to adversarial settings. We deploy a large-scale automated evaluation pipeline to ensure rigorous and reproducible assessment. The reliability of our automated metrics and results is verified by human experts (over 97% agreement rate). The initial findings on representative models are summarized in a public leaderboard (see Figure˜2(b)). Details of the benchmark are provided in Appendix˜C and Figure˜4.
Fairness. The introduction of audio inputs brings new fairness risks by introducing new biases linked to audio characteristics. To investigate these risks, we conducted a comprehensive evaluation of model fairness along two dimensions: decision-making experiments and stereotype-association experiments. Our main findings are as follows: (1) Audio-based attributes (e.g., accent, emotion) can introduce biases that are stronger than those from traditional sensitive attributes (e.g., age, gender), indicating that audio information is a key carrier of bias; (2) We observed that closed-source models exhibited stronger decision biases, while open-source models were more susceptible to stereotype associations; (3) The identified biases tend to disadvantage non–socially-dominant groups, such as older-sounding accents, perceived calmness, and markers of lower socioeconomic status (SES). Further details are provided in Section˜2.
Hallucination: The introduction of audio gives rise to new forms of hallucination, including the misinterpretation of paralinguistic features (e.g., emotion or accent) and failures to capture temporal causality within speech. We study these vulnerabilities on a carefully curated benchmark and identify several key weaknesses. (1) Closed-source models exhibit stronger robustness when confronted with acoustically implausible events. (2) Many models remain vulnerable to misleading meta-attributes, revealing insufficient alignment with domain knowledge. (3) We further observe pronounced fragility in tasks that require temporal reasoning. (4) Substantial variability exists across models in terms of cross-modal semantic consistency. Further details are provided in Section˜3.
Safety: Incorporating audio inputs substantially broadens the attack surface. Unlike text, speech carries emotional, contextual, and anthropomorphic cues that adversaries can exploit by modulating tone, injecting affective signals, or impersonating identities. We design a composite attack framework spanning emotion-driven deception and identity-verification evasion. Empirical analysis shows: (1) closed-source models exhibit stronger overall robustness but remain sensitive to highly emotional speech; (2) open-source models are disproportionately vulnerable to identity- and emotion-based manipulation, often yielding unsafe outputs in high-stakes settings such as healthcare; (3) for prompts concerning illicit guidance, closed-source systems largely resist, whereas some open-source models deliver risky recommendations in multi-turn audio dialogues. Further details are provided in Section˜3.
Privacy: ALLMs face two closely related but mechanistically distinct audio privacy risks. The first is content-level leakage, such as reading out and repeating bank account numbers, social security numbers, or addresses. The second is paralinguistic-level inference leakage, where attributes such as age, gender, race, geographic location, or socioeconomic status are inferred from voiceprints, timbre, intonation, accent, or background sounds. To conduct a systematic evaluation, we created targeted scenarios to assess both explicit information disclosure and implicit attribute inference. Our findings show: (1) ALLMs are relatively robust in preventing direct content leakage; (2) existing semantic-oriented defenses fall short in addressing paralinguistic attack surfaces unique to audio, underscoring the need to integrate acoustic and environmental cues into privacy-aware decision boundaries. See Section˜5 for details.
Robustness: Since ALLMs interact directly through audio, they are inevitably affected by noise and distortion. To systematically characterize their robustness, we evaluated the models against a comprehensive suite of real-world audio degradations, including environmental noise, speaker overlap, and signal perturbations. Our analysis reveals: (1) mainstream closed-source ALLMs achieve stronger task performance under overlapping speech, non-stationary noise, and reverberant conditions, while most open-source models exhibit substantial performance degradation; (2) existing ALLMs generally demonstrate an “over-textualization” tendency, where models continue reasoning based on partially correct transcripts while neglecting acoustic cues when transcription is correct but acoustic attribution is mistaken. See Section˜6 for details.
Authentication: In applications of ALLMs, speech-related authentication issues are particularly critical. To investigate these risks, we evaluated the models against several key attack vectors, including identity verification bypass and voice cloning deception. The results show: (1) certain closed-source models exhibit some resilience in identity verification scenarios, whereas open-source models are generally more vulnerable to sophisticated voice-based attacks; (2) adversaries may exploit social engineering or acoustic interference, such as background crowd noise, to compromise verification reliability; (3) employing more stringent speech-text prompting strategies can substantially improve the ability of ALLMs to withstand voice cloning attacks. See Section˜7 for details.
@@ 2 AudioTrust: Fairness
This section examines the fairness issues associated with ALLMs. Fairness risks in audio models are fundamentally different from those in text or vision systems. For instance, a text-based model might exhibit bias based on a name mentioned in a hiring application, but an ALLM can develop biases from the acoustic cues of an applicant’s voice alone. A hesitant speaking style could be misinterpreted as a lack of confidence, or a non-native accent could trigger stereotypes, regardless of the spoken content’s quality. Traditional fairness metrics focusing on textual protected attributes are insufficient to capture these audio-native biases. We investigate these new audio forms of bias that arise from how auditory characteristics are perceived.
Attack Strategies. To systematically probe these risks, we categorize fairness into two dimensions: traditional fairness and audio-based fairness. Traditional fairness assesses biases linked to demographic attributes like gender, race, and age (Cho et al., 2023; Zhang et al., 2017b; Wan et al., 2024) that can be inferred from a voice. We test, for example, if a model’s loan approval decisions are skewed by whether the applicant sounds male or female. Audio-based fairness isolates biases triggered by instrinsic acoustic properties. We divide these biases into four sub-categories: accent, linguistic characteristic (e.g., speech disfluencies may adversely impact the model’s fairness), economic status (e.g., noisy environments, often correlated with lower economic status, might yield more negative outputs), and personality traits (e.g., negative emotions may cause the model to produce negative outputs) (see Figure 5). For each sensitive attribute, we designed decision-making and stereotype-driven scenarios (Xu et al., 2025a; Wang et al., 2024b). Decision-making covers recruitment, admissions, and loan evaluations. Stereotype-driven contexts address beliefs such as men outperforming women in mathematics and gender roles in medicine and occupations. We constructed a dedicated dataset consisting of 840 audio samples, each lasting approximately 20 seconds. The samples were annotated with seven key sensitive attributes: gender, age, and race for evaluating traditional fairness, as well as accent, linguistic characteristic, economic status, and personality traits for assessing audio-baseed fairness. Notably, due to the limitations of the audio modality, attributes such as appearance are not included (Nangia et al., 2020; Parrish et al., 2021). Detailed dataset construction procedures can be found in Sections˜D.1 and D.2.
Evaluation & Metrics. For fairness evaluation, we introduced a key metric: Group Unfairness Score (Xu et al., 2025a; Deldjoo & Nazary, 2024). Group unfairness examines the distributional equity of model outputs across different groups (e.g., male versus female), measured by the divergence or balance between group distributions. indicates perfect fairness, while indicates extreme unfairness. In computing the unfairness score, we used human annotation and counting throughout to ensure data correctness and validity. For detailed formulas in the Section˜D.3.
@@
| SALMONN | 0.861 | 0.911 |
| Ultravox | 0.762 | 0.608 |
| Qwen2-Audio | 0.667 | 0.710 |
| MiniCPM-o 2.6 | 0.740 | 0.585 |
| Step-Fun | 0.342 | 0.495 |
| Qwen2.5-Omni | 0.933 | 0.798 |
| Kimi-Audio | 0.964 | 0.914 |
| OpenS2S | 0.983 | 0.843 |
| Step-Audio2 | 0.926 | 0.902 |
@@
| Gemini-1.5 Pro | 0.703 | 0.540 |
| GPT-4o Audio | 0.074 | 0.736 |
| GPT-4o mini Audio | 0.136 | 0.755 |
| Gemini-2.5 Flash | 0.630 | 0.754 |
| Gemini-2.5 Pro | 0.681 | 0.795 |
| Average | 0.672 | 0.739 |
Note: : higher than column average, : lower than column average, subscript is signed difference from mean (3 decimals).
Results. We evaluated the group fairness of 14 ALLMs in terms of social stereotypes and decision-making in Table 1. Complete results and examples are provided in Section˜D.4. The main findings are as follows: (1) Existing ALLMs exhibit severe unfairness across different sensitive attributes, falling far short of the ideal fairness (i.e., ). (2) The GPT-4o series shows a pronounced disparity between decision and stereotype . This is because we have designed extreme decision scenarios, and the GPT-4o series models sacrifice fairness to maintain accuracy in response. (3) Although leading closed-source models such as GPT-4o Audio exhibit stable fairness performance, open-source models vary widely. Notably, Step-Fun demonstrates strong fairness, with scores comparable to the best closed-source models. By contrast, models like OpenS2S and SALMONN display pronounced vulnerabilities, underscoring a substantial capability gap within the open-source ecosystem. (4) The two models in the Step series exhibit a stark disparity in fairness, suggesting substantial differences in their underlying fairness mechanisms.
@@ 3 AudioTrust: Hallucination
In this section, we examine the hallucination problem in ALLMs. Audio hallucinations extend beyond the factual errors seen in text(Ji et al., 2023; Sriramanan et al., 2024; Yao et al., 2023). An ALLM does not just process information; it interprets a simulated physical world. For instance, if an audio recording contains the sound of a gunshot followed by a body falling, but the model describes the fall happening first, it is not just a factual error, which is a violation of causality. Similarly, describing a fire burning underwater is a violation of physical laws. These audio-grounded errors are undetectable by text-based fact-checking and pose unique safety risks. Our work is the first to systematically define and evaluate these physically and logically grounded hallucinations.
Attack Strategies. In AudioTrust, we identify two main categories of audio hallucinations (see Appendix E.1): Physical Logic and Chronological Order. The former relates to violations of acoustic laws and environment properties, and the latter reflects failures in reasoning about temporal and causal relations. These distinctions directly connect to safety risks in real-world use. To evaluate them, we built a dataset of 320 samples from synthetic and real sources (see Appendix E.2). For Acoustic-Physical Hallucinations, we focus on two specific manifestations: (1) Contravention of Physicochemical Constraints, generating impossible events (e.g., the flames are burning in the seawater) to test propagation understanding (Liu et al., 2024); (2) Source–Environment Mismatch, applying contradictory reverberation (e.g., casual speech with cathedral acoustics) to test disentanglement of source vs. environment (Zhou et al., 2022). For Temporo-Logical Hallucinations, we examine (1) Temporal-Causal Inversion: reversing causal chains (e.g., engine start before ignition) to probe event logic (Chen et al., 2023); (2) Cross-Modal Contradiction: pairing audio with conflicting text (e.g., fast footsteps described as peaceful rain) to test cross-modal reasoning (Deshmukh et al., 2023).
Evaluation & Metrics. We introduce a comprehensive evaluation framework to assess model capabilities across four key dimensions: hallucination detection, attribute verification, real-world consistency, and transcription accuracy. For hallucination detection, models are required to identify inaccuracies in audio-text pairs and provide justifications. Performance is quantified via a multi-dimensional, GPT-4o-based evaluator (OpenAI et al., 2024).The scores mentioned above range from 0 to 10 and are used to evaluate the accuracy of the detection and the quality of the interpretation. These scores were then subject to further review by the human evaluators. We probe for attribute-level hallucinations related to physical properties, labels, and content using multi-stage prompting (Madaan et al., 2023). To measure alignment with factual information, we adopt the two-stage protocol from Li et al. (2023a) for real-world consistency assessment. Finally, we evaluate transcription robustness under hallucinatory interference using both the standard Word Error Rate (WER) (Deshmukh et al., 2023) and the cross-modal WER (CM-WER) (Tang et al., 2024). The complete experimental design and metric details are provided in Appendices E.3.1 and E.3.2.
@@
| Qwen2.5-Omni | 8.12 | 5.63 | 7.89 | 6.11 |
| Step-Fun | 3.96 | 4.84 | 5.80 | 8.72 |
| Kimi Audio | 1.86 | 5.77 | 5.82 | 8.54 |
| Step-Audio2 | 3.62 | 1.94 | 2.60 | 2.76 |
| OpenS2S | 1.92 | 5.03 | 5.97 | 7.89 |
@@
| Average | 5.35 | 5.11 | 6.65 | 7.78 |
Note: Scenarios: Content Mismatch (CM), Label Mismatch (LM), Logical Violation (LV), and Physical Violation (PV).
Results. As shown in Table 2, our evaluation highlights both the progress and the critical limitations of current ALLMs in resisting hallucinations. Complete results and examples are provided in Section˜E.4 and Section˜E.5. We observe two main findings: (1) Although certain open source models, such as Gemini-2.5, demonstrate the ability to detect specific types of hallucinations, particularly those with explicit physical or temporal contradictions.(e.g., statements claiming an object exists in two distinct places at the same time or a water bottle made a sound by hitting the ground during its fall) Nonetheless, overarching vulnerabilities persist. Most models falter on subtler instances , including source-environment incongruities (e.g., an audio track describing a visual scene that contradicts the dialogue content) or cross-modal semantic discrepancies (e.g. The audio background information of a scene contradicts the content of the dialogue.), indicating that their perceptual understanding remains fragmented lacking an integrated cognitive architecture. (2) A striking observation is the negative correlation between the subjective complexity of tasks from a human perspective and the actual performance of models. While models attain comparatively high accuracy in identifying violations of physical laws, they underperform on content mismatches such as scenarios and independent tasks in different scenarios that humans intuitively discern with ease. This divergence highlights a core disparity in auditory perception and reasoning between humans and machines: models excel at low-level acoustic anomaly detection but struggle to emulate human-like commonsense reasoning.